The number of internal disk drives (in milions) made at a plant in Talwan during the past 5 years follows: Disk Drives 142 156 196 204 216 Year a) Using simple linear regression, the forecast for the number of disk drives to be made next year disk drives (round your response to one decimal place) b) The mean squared error(MSE) when using simple inear c) The mean absolute percentage error (MAPE) when using simple linear regression . regressiondrives (round your response to one decimal place) % round your response to one decimal place
Answer #1

A.

Annual No. of Disk Drives 250T y 19.6x+124 Re* E 150 100 50 Year

According to the above graph of trend line and regression equation,

No. of disk drives =124 + 19.6*year

The data is of past 5 year, and present year is 6th year. So, next year is 7th year.

Forecast of No. of disk drives in year 7 =124 + 19.6*7 = 261.2 or 261

B.

Year Disk Drives Regression forecast Absolute error Square of absolute error
1 142 143.6 1.6 2.56
2 156 163.2 7.2 51.84
3 196 182.8 13.2 174.24
4 204 202.4 1.6 2.56
5 216 222 6 36
MSE = 53.44

MSE = mean of the square error = (2.56+51.84+174.24+2.56+36)/5 =  53.44

C.

Year Disk Drives Regression forecast Absolute error Absolute percentage error (%)
1 142 143.6 1.6 1.13
2 156 163.2 7.2 4.62
3 196 182.8 13.2 6.73
4 204 202.4 1.6 0.78
5 216 222 6 2.78
MAPE = 3.21

MAPE = Mean of the absolute percentage error = (1.13+4.62+6.73+.78+2.78)/5 =  3.21%